OpenStudy (anonymous):

Mike shoots a large marble (Marble A, mass: 0.05 kg) at a smaller marble (Marble B, mass: 0.03 kg) that is sitting still. Marble A was initially moving at a velocity of 0.6 m/s, but after the collision it has a velocity of –0.2 m/s. What is the resulting velocity of marble B after the collision? Be sure to show your work for solving this problem along with the final answer.

OpenStudy (anonymous):

0.4 m/s

OpenStudy (anonymous):

Explain please.

OpenStudy (anonymous):

Step by step.

OpenStudy (anonymous):

Cant...its all in my insane head...im just pretty sure thats correct...

OpenStudy (anonymous):

I can't just put down 0.4 m/s.

OpenStudy (anonymous):

Please . . .

OpenStudy (anonymous):

@lolrocker1

OpenStudy (anonymous):

Dude...if i knew id explain it...but i dont..that seems logicly what is correct to me...i think i got it because the decelerati...wtf i dont know man...im probably wrong...

OpenStudy (anonymous):

@lolrocker1 I found the correct answer I just thought I'd let you know :) To find the Momentum of Marble A multiply the Mass(0.05) and the velocity (0.6) 0.05 x 0.6 = 0.03kg m/s. To find the Momentum of Marble B multiply the mass(0.03) and the velocity(0). 0.03 x 0 = 0. To find the velocity of Marble A Multiply 0.05 x -0.2. Simplify: 0.03 + 0 = 0.05 x -0.2 + 0.03x 0.03 + 0 = -0.01 + 0.03x 0.03 = -0.01 + 0.03x Subtract -0.01 to both sides: 0.04 = 0.03x Divide: 0.04 / 0.03 ≈ 1.3 So the velocity of Marble B after the Collison is 0.04 / 0.03 or ≈ 1.3.